# Analytic Continual Test-Time Adaptation for Multi-Modality Corruption

### v1

@@ Abstract
Test-Time Adaptation (TTA) aims to help pre-trained model bridge the gap between source and target datasets using only the pre-trained model and unlabelled test data. A key objective of TTA is to address domain shifts in test data caused by corruption, such as weather changes, noise, or sensor malfunctions. Multi-Modal Continual Test-Time Adaptation (MM-CTTA), an extension of TTA with better real-world applications, further allows pre-trained models to handle multi-modal inputs and adapt to continuously-changing target domains. MM-CTTA typically faces challenges including error accumulation, catastrophic forgetting, and reliability bias, with few existing approaches effectively addressing these issues in multi-modal corruption scenarios. In this paper, we propose a novel approach, Multi-modality Dynamic Analytic Adapter (MDAA), for MM-CTTA tasks. We innovatively introduce analytic learning into TTA, using the Analytic Classifiers (ACs) to prevent model forgetting. Additionally, we develop Dynamic Selection Mechanism (DSM) and Soft Pseudo-label Strategy (SPS), which enable MDAA to dynamically filter reliable samples and integrate information from different modalities. Extensive experiments demonstrate that MDAA achieves state-of-the-art performance on MM-CTTA tasks while ensuring reliable model adaptation.
1 Introduction
@@
Test-Time Adaptation (TTA) aims to help the pre-trained model bridge the gap between the source domain and the target domain (Wang et al., 2021; Liang et al., 2024). Unlike Unsupervised Domain Adaptation (UDA) (Zhang et al., 2015; Liang et al., 2024), TTA performs adaptation without the need for any source data (i.e., pre-trained dataset), which not only saves computational resources by avoiding retraining but also preserves the privacy of the source data. One key TTA application is addressing the problem of domain shift from source data to corrupted test data, where the corruption is often caused by external factors (e.g., weather changes, ambient noise) or sensor malfunctions. As an extension of TTA, Continual Test-Time Adaptation (CTTA) has been proposed to align with real-world scenarios where domain shifts usually are dynamic (Wang et al., 2022). Challenges in CTTA mainly consist of error accumulation and catastrophic forgetting. Error accumulation, stemming from incorrect pseudo-labels, can mislead models’ adaptation and potentially lead to collapse (Chen et al., 2019). Catastrophic forgetting refers to the loss of knowledge from the source data during continuous adaptation, reducing the model’s generalization ability (McCloskey & Cohen, 1989). To address these challenges, various CTTA methods have been proposed, yielding promising results in corruption-related tasks (Wang et al., 2022; Gao et al., 2023; Niu et al., 2022).
Most existing CTTA approaches focus solely on single-modal scenarios, paying less attention to multi-modal applications. Compared with TTA and CTTA, Multi-Modal Continual Test-Time Adaptation (MM-CTTA) (Cao et al., 2023) shows greater potential for real-world applications, as multi-modal data integrates a broader range of information than single-modality adaptation, resulting in more robust networks (Radford et al., 2021). However, applying existing CTTA methods to MM-TTA by simply replacing the backbone with muiti-modal encoders is less optimal. MM-CTTA performance can easily suffer from reliability bias, where intra-modality domain shifts increase information discrepancies in downstream fusion networks (Yang et al., 2024). Such effect becomes more pronounced when modality corruption changes dynamically. Although a few works address Multi-Modal Test-Time Adaptation (MM-TTA) (Yang et al., 2024; Lei & Pernkopf, 2024), they struggle with model forgetting during continuous adaptation and the challenges posed by interleaved modality corruption.
To further illustrate, we use a representative TTA method, TENT (Wang et al., 2021), as an example in Fig.1. The results indicate that traditional TTA methods face significant limitations due to error accumulation, catastrophic forgetting, and reliability bias. These challenges suggest that methods designed for MM-CTTA must meet three key requirements: (1) effectively mitigating error accumulation during adaptation, (2) retaining knowledge of the source data after adapting to various domains, and (3) dynamically suppressing the influence of corrupted modalities while prioritizing more reliable ones. To the best of our knowledge, the method closest to meeting these criteria is CoMAC (Cao et al., 2023), which focuses on segmentation tasks in living environments rather than addressing modality corruption.
In this paper, we propose a new approach named the Multi-modality Dynamic Analytic Adaptor (MDAA) to address the challenges in MM-CTTA. MDAA comprises three primary components: (i) the Analytic Classifiers (ACs), (ii) the Dynamic Selection Mechanism (DSM), and (iii) the Soft Pseudo-label Strategy (SPS). ACs update the model by addressing a recursive ridge regression problem, optimizing on both new target data and learned knowledge to avoid catastrophic forgetting. DSM selectively updates each AC based on its output reliability, thereby alleviating reliability bias. The SPS enhances the model’s robustness to label noise by assigning varying probabilities to multiple labels, which mitigates potential error accumulation. contribution of thiss work can be summarised as follows:
1. We c
1). We propose a method, MDAA, for a more challenging TTA task named MM-CTTA, and explain why typical TTA methods are not well-suited for this task, as illustrated by a example in Fig.1.
2). We innovatively apply AC to TTA to keep the model from catastrophic forgetting during adaption. We propose DSM and SPS to further dynamically integrate the features of different modal thus mitigating reliability bias the error accumulation.
3). We design two MM-CTTA tasks to meet the needs of real-world environments. Extensive experiments demonstrate that MDAA achieves SOTA performance, surpassing previous methods by up to 6.22% and 6.84% in the two tasks respectively.
2 Related Works
2.1 Test-Time Adaptation
@@
Test-Time Adaptation (TTA) focuses on enabling a pre-trained model to adapt to a new target domain without requiring access to the source domain data used to initially train the model. A major challenge in TTA is error accumulation, which affects methods that rely on pseudo-labeling; incorrect predictions can mislead the adaptation process, potentially causing the model to collapse (Chen et al., 2019). One of the early solutions, TENT, addresses this by only updating the model’s batch normalization (BN) layers through entropy minimization (Wang et al., 2021). Subsequent research (Niu et al., 2023; Gong et al., 2022a; Zhou et al., 2023) has further mitigated this issue by filtering out low-confidence predictions using carefully designed thresholds and updating the layer normalization (LN) layers for more robust performance.
2.2 Continual Test-Time Adaptation
Continual Test-Time Adaptation (CTTA) extends TTA to scenarios where the target domain changes continuously in an online manner without access to source data. The additional challenge CTTA faces is known as catastrophic forgetting, which occurs when a model adapts to the target domain, leading to the loss of knowledge acquired from the source domain and dimension of generalization ability. To solve this problem, some studies turn to Continual Learning (CL) and achieve great success (Niu et al., 2022; Cao et al., 2023). We follow this trend and introduce a novel CL approach called Analytic Continual Learning (ACL) in MDAA.
ACL provides a global optimal solution through matrix inverse operations (Guo & Lyu, 2004). To address the out-of-memory issue caused by large inverse matrices, Zhuang et al. (2021a) demonstrated that iterative computation using block-wise data achieves results equivalent to joint computation, making analytic learning highly effective in continual learning. By treating data from different time periods as blocks, ACL allows for recursive computation, with the final result being as accurate as if all data were processed simultaneously. Thanks to its non-forgetting properties, ACL has shown strong performance across various CL tasks in recent years (Zhuang et al., 2022; 2023). Inspired by ACL’s success in CL tasks, we apply ACL to TTA for the first time in this work, implementing several enhancements to address the issue of catastrophic forgetting.
2.3 Multi-modality Test-Time Adaptation
@@
Multi-modality Test-Time Adaptation (MM-TTA) seeks to enhance model reliability by incorporating multi-modal data into the TTA task. However, imbalances in inter-modal reliability can result in significant performance degradation, known as reliability bias (Wang et al., 2020). Most existing MM-TTA models address this issue by independently updating BN or LN layers of each feature encoder, followed by a weighted fusion mechanisms (Shin et al., 2022; Cao et al., 2023; Xiong et al., 2024). Although this allows more reliable modalities to carry more weight during fusion, the approach remains relatively shallow in terms of information integration.
@@
A recent model called READ (Yang et al., 2024) introduced a more advanced approach by fusing features through a Vision Transformer (ViT) (Dosovitskiy et al., 2020) block, which allows for the preservation of parameters inherited from source data while effectively integrating inter-modal information (Vaswani, 2017; Gong et al., 2022b). READ achieves reliable adaptation by modulating only the fusion layer within the attention module of the ViT block. Follow-up work (Lei & Pernkopf, 2024) aimed to improve performance by updating both the feature encoders and the fusion layer. However, this approach requires prior knowledge about which modalities are corrupted, making it less capable for real-world scenario.
In this paper, our MDAA approach introduces classifiers for each feature encoder and the fusion network. By adjusting the parameters of each classifier individually, MDAA maximizes the use of information from each modality, thus mitigating the effects of reliability bias. Crucially, these classifier updates for upstream and downstream blocks do not cause conflicts, as the entire pre-trained model remains frozen throughout the process. This property enables MDAA to adapt to changing modality corruption without extra prior knowledge, distinguishing it from existing MM-TTA approaches.
3 Method
In this section, we first define the challenging MM-CTTA setting and introduce notations for key concepts in Sec. 3.1. We then introduce the proposed Multi-modality Dynamic Analytic Adapter (MDAA) approach, which includes Analytic Classifiers (ACs) integrated with pre-trained multi-modal encoders (Sec. 3.2), the Dynamic Selection Mechanism (DSM) (Sec. 3.3), and the Soft Pseudo-label Strategy (SPS) (Sec. 3.4). An overview of MDAA is illustrated in Fig. 2, and the pseudo-code for MDAA is provided in Appendix B.
3.1 Problem Definition and Notations
@@
In this paper, we focus on an audio-video classification task as an example, using two modalities for illustration without loss of generality. In MM-CTTA, the pre-trained model is trained on a labeled source dataset in source domain S, where and represent the audio and video training data matrices respectively. represents the corresponding one-hot label set. During the adaptation phase, for each timestamp in the target domain T, the model must perform inference and update its parameters based on the unlabeled test dataset . The suffix indicates the current target domain, as it shifts continually. Note that only the test dataset at timestamp is available for updating the parameters .
3.2 Source Model and Analytic Classifier Configuration
In the context of multi-modality classification task, we propose to integrate the ACs as classifiers into a typical extraction-fusion approach. The standard structure can be represented as follows:
where and represent upstream feature encoders for audio and video modalities, respectively, indicates the fusion operation implemented as concatenation in this paper, denotes the downstream fusion network and indicates the classifier.
Specifically, we leverage multiple ACs as classifiers for extracted features of each modality and the fused features, with each classifier making an independent prediction. Each AC is of the same structure as a two-layer fully connected network, denoted by . Features are first non-linearly projected into a higher dimensional space by the feature expansion layer , enhancing their expressiveness (Zhuang et al., 2022; 2021b). The projected features are then passed through the linear layer for classification. The feature expansion layer remains frozen, while the linear layer requires updates during the adaptation process. For simplicity, we combine the feature encoder and the feature expansion layer into a single function, denoted as , and refer to the projected features as . Consequently, MDAA yields three classifiers as follows:
Since all ACs can be represented by the same equation, we will not distinguish between them except for special needs in the following discussion. Following previous ACL works Zhuang et al. (2022), the classifier are updated by solving the ridge regression to optimize on the source data . To further solve the class imbalance problem within the source data, inspired by Fang et al. (2024), we formulate the optimization problem as follows:
@@ 3.2 Source Model and Analytic Classifier Configuration
where is the number of classes in and is the number of samples in from category to which sample belongs. Following the ridge regression solution, the solution to optimization problem is given in Theorem 1.
Theorem 1. The optimal solution to Formula 4 is given as
@@
where and . The proof of Theorem 1 is provided in Appendix A. In addition to the classifier weights , a memory bank needs to be constructed during the training phase. Unlike other methods (Cao et al., 2023; Zhang et al., 2023; Xiong et al., 2024), our memory bank contains only two types of matrices, which can be represented as , where
@@
Both and are used to extract and preserve the learned knowledge from the source dataset, which cannot be accessed during adaptation. Therefore weight can be further rewritten as
3.3 Dynamic Selection Mechanism
When adapting to sequentially incoming multi-modal data in MM-CTTA, some modality may be unreliable due to corruption. Such corrupted data can mislead the model to learn incorrect information. To address the reliability bias, we propose Dynamic Selection Mechanism (DSM) to determine whether each AC should be updated in a dynamic way. M first identifies the most reliable classifier among the three (i.e., a, v, m) as the leader, while the other classifiers are treated as follower.
Specifically, the leader is determined by comparing the maximum probability from each classifier’s distribution. For each sample, the model prediction corresponds to the leader’s prediction, and the leader will not be updated to maintain class balance. Whether to update each follower in the model depends on the comparison between its maximum probability distribution and the leader’s maximum probability distribution . We consider four possible scenarios, as illustrated in Fig.3:
(i). Close Distributions: and are quite close and refer to the same label. In this case, the follower is not updated, as this would only reinforce what it has already learned, potentially leading to an imbalanced class distribution.
(ii). Different Labels with Close Probabilities: and are quite close while they refer to different labels. The follower is also not updated in this scenario because there is no certainty that the leader’s result is correct, and updating may introduce errors.
(iii). Evenly Distributed Probabilities: Both the leader and follower have evenly distributed probabilities with no significant difference between labels. Again, no update occurs.
(iv). Significant Difference: The leader has a higher probability label, while the follower has a more even probability distribution. In this scenario, the follower should be updated.
In general, cases (i),(ii) and (iii) all belong to a small gap between and , while case (iv) belongs to a larger gap. Therefore, given a pre-defined threshold , DSM can be noted as:
3.4 Soft Pseudo-label Strategy
Using soft labels in self-supervised learning is quite popular as it preserves a broader range of possible outcomes compared to hard (one-hot) label learning (Müller et al., 2019; Hinton et al., 2015). Inspired by such trend, we use soft pseudo-labels to update the ACs during MM-CTTA, as illustrated in Fig.2(C). For each test sample , we choose the top classes of leader’s distribution set and assign them with weights , () respectively. The reconstructed label for sample through SPS can be represented as
Since ACL considers global optimization, which means it accounts not only for the input data at timestamp , but also all previous data processed by the model, including and . With the reconstructed label throughout , the optimization problem for weight matrix can be represented as:
It is important to note that while we treat source data and target data separately, we adjust the category balance for the source dataset by assigning weights to each class. In contrast, it is challenging to apply this approach to the target dataset, as it is impossible to know the exact number of samples in each category within the target domain. However, with the help of the DSM, category balance can still be maintained, as unimportant samples have been filtered out. Given that the weights added to the sample data are expected to average out to 1, the impact of each sample from both the source and target domains on the model can be considered equivalent.

### v2

@@ Abstract.
Test-Time Adaptation (TTA) enables pre-trained models to bridge the gap between source and target datasets using unlabeled test data, addressing domain shifts caused by corruptions like weather changes, noise, or sensor malfunctions in test time. Multi-Modal Continual Test-Time Adaptation (MM-CTTA), as an extension of standard TTA, further allows models to handle multi-modal inputs and adapt to continuously evolving target domains. However, MM-CTTA faces critical challenges such as catastrophic forgetting and reliability bias, which are rarely addressed effectively under multi-modal corruption scenarios. In this paper, we propose a novel approach, Multi-modality Dynamic Analytic Adapter (MDAA), to tackle MM-CTTA tasks. MDAA introduces analytic learning—a closed-form training technique—through Analytic Classifiers (ACs) to mitigate catastrophic forgetting. Furthermore, we design the Dynamic Late Fusion Mechanism (DLFM) to dynamically select and integrate reliable information from different modalities. Extensive experiments show that MDAA achieves state-of-the-art performance across the proposed tasks.
1. Introduction
@@
Test-Time Adaptation (TTA) aims to enable pre-trained models to bridge the gap between the source domain and the target domain (Wang et al., 2021; Liang et al., 2024). Unlike Unsupervised Domain Adaptation (UDA) (Zhang et al., 2015), TTA performs adaptation without any source data (i.e., pre-trained dataset), which not only saves computational resources by avoiding retraining but also preserves the privacy of the source data. One key application of TTA is to address the problem of domain shift from source data to corrupted test data, where the corruption is often caused by weather changes, ambient noise or sensor malfunctions. In these scenarios, Multi-Modal Test-Time Adaptation (MM-TTA) offers a distinct advantage over mono-modal TTA, as domain shifts often affect only a subset of modalities, while others remain relatively stable and informative (Radford et al., 2021; Yang et al., 2024). This allows the model to leverage uncorrupted modalities to guide adaptation, thereby enhancing robustness in complex real-world environments.
To better reflect real-world settings where distribution shifts are continuous and multi-faceted, Multi-Modal Continual Test-Time Adaptation (MM-CTTA) has been proposed (Cao et al., 2023). MM-CTTA addresses more complicated domain shifts, where the target distribution evolves both over time and across modalities. This setting is common in real-world scenarios, particularly for general-purpose AI agents (e.g., autonomous driving, robotic systems) that rely on multiple sensors for perception, as illustrated in Figure 1.A. However, achieving reliable MM-CTTA remains challenging, mainly due to two key issues. The first is reliability bias, which arises from intra-modality domain shifts and exacerbates inconsistencies within downstream fusion networks (Yang et al., 2024). The second is catastrophic forgetting, where knowledge acquired from the source domain deteriorates during continual adaptation, ultimately reducing the model’s generalization capability (McCloskey and Cohen, 1989).
MM-CTTA models must continuously adapt to multi-modal domain shifts while mitigating modality collapse and ensuring reliability in real-world scenarios. While recent researches on TTA have achieving promising results of in catastrophic forgetting (Wang et al., 2022; Gao et al., 2023; Niu et al., 2022) and reliability bias (Yang et al., 2024; Lei and Pernkopf, 2024) respectively, directly combining CTTA methods with MM-TTA works often falls short. This is because dynamically changing corruption patterns magnify reliability bias and create additional complexities, as corruption can unpredictably switch between modalities during continual adaptation. Methods explicitly designed for MM-CTTA remain scarce. The closest existing approach, CoMAC (Cao et al., 2023), focuses primarily on segmentation tasks in dynamic environments rather than addressing modality corruption.
In this paper, we propose a new approach named Multi-modality Dynamic Analytic Adaptor (MDAA) to address the challenges in MM-CTTA. MDAA mainly comprises two primary components: (i) the Analytic Classifiers (ACs), and (ii) the Dynamic Late Fusion Mechanism (DLFM). The ACs update the model by addressing a recursive ridge regression problem, optimizing on both new target data and learned knowledge to avoid catastrophic forgetting, such approach have already been used in continual learning (CL) field (Zhuang et al., 2022, 2023). Moreover, each feature type in the multi-modal model is configured with an individual AC in our framework. As a result, MDAA can derive predictions not only from the downstream fusion network but also from the feature encoders, marking a significant departure from conventional multi-modal paradigms.
To effectively utilize outputs from different ACs, we proposed a tailored late fusion strategy, DLFM, which compensates for and integrates predictions across modalities. DLFM selects reliable predictions from modality-specific ACs and propagates them to guide the updates of ACs associated with corrupted modalities, thereby alleviating reliability bias. To rigorously assess the robustness and adaptability of MDAA, we design two challenging evaluation settings: (1) adaptation over extremely long phases and (2) scenarios with alternating modality corruption. These settings provide a rigorous test of MDAA’s robustness and adaptability under real-world conditions. The key contributions of this work are listed as follows:
1). We introduce ACs to TTA into the TTA framework to prevent the model from catastrophic forgetting. We propose DLFM to adaptively fuse features from multiple modalities, thereby mitigating reliability bias.
2). We design two MM-CTTA benchmarks featuring modality corruption switching and prolonged adaptation, better capturing real-world dynamics, pushing the boundaries of MM-CTTA.
3). MDAA achieves state-of-the-art performance, outperforming prior methods by up to 6.22% and 6.84% in two benchmarks, while being more efficient in runtime and memory usage.
2. Related Works
2.1. Test-Time Adaptation
@@
TTA focuses on enabling a pre-trained model to adapt to a new target domain without requiring access to the source domain data. However, simply relying on incorrect predictions during adaptation can mislead the update process, potentially leading to model collapse (Chen et al., 2019). TENT, as one of the first methods solve such issue, addresses this by only updating the model’s batch normalization (BN) layers through entropy minimization (Wang et al., 2021). Subsequent research (Niu et al., 2023; Gong et al., 2022a; Zhou et al., 2023) has further mitigated this issue by filtering out low-confidence predictions using carefully designed thresholds and updating the layer normalization (LN) layers for more robust performance.
2.2. Multi-modality Test-Time Adaptation
@@
MM-TTA (Zhao et al., 2025; Guo and Jin, 2025; Yang et al., 2024; Wang et al., 2020) seeks to enhance model reliability by incorporating multi-modal data into the TTA task. However, imbalances in inter-modal reliability can result in significant performance degradation, known as reliability bias (Wang et al., 2020), shown in Figure 1.B. Existing MM-TTA models address this issue by independently updating BN or LN layers of each feature encoder, followed by a weighted fusion mechanisms (Shin et al., 2022; Cao et al., 2023; Xiong et al., 2024). Although this allows more reliable modalities to carry more weight during fusion, the approach remains relatively shallow in terms of information integration.
@@
A recent model called READ (Yang et al., 2024) introduced a more advanced approach by fusing features through a Vision Transformer (ViT) (Dosovitskiy et al., 2020) block, which allows for the preservation of parameters inherited from source data while effectively integrating inter-modal information (Vaswani, 2017; Gong et al., 2022b). READ achieves reliable adaptation by modulating only the fusion layer within the attention module of the ViT block. Follow-up work (Lei and Pernkopf, 2024) aimed to improve performance by updating both the feature encoders and the fusion layer. However, this approach requires prior knowledge about which modalities are corrupted, making it less capable for real-world scenario.
2.3. Analytic Continual Learning
When extending TTA to a continual setting, a significant challenge that emerges is catastrophic forgetting—a phenomenon where adaptation to the target domain leads to degradation of knowledge learned from the source domain and compromises generalization (as illustrated in Figure 1.C). Recent works have explored the integration of Continual Learning (CL) into TTA, demonstrating promising results (Niu et al., 2022; Cao et al., 2023). Following this line of research, we incorporate a novel CL paradigm, Analytic Continual Learning (ACL), into our proposed MDAA framework.
ACL provides a global optimal solution through matrix inverse operations (Guo and Lyu, 2004). To address the out-of-memory issue caused by large inverse matrices, (Zhuang et al., 2021) demonstrated that iterative computation using block-wise data achieves results equivalent to joint computation, making analytic learning highly effective in continual learning. By treating data from different time periods as blocks, ACL allows for recursive computation, with the final result being as accurate as if all data were processed simultaneously. Thanks to its non-forgetting properties, ACL has shown strong performance across various CL tasks in recent years (Zhuang et al., 2022, 2023).
3. Multi-modality Dynamic Analytic Adapter
In this section, we first formalize the MM-CTTA problem and introduce the key notations in Sec.3.1. Then present the proposed MDAA framework, which incorporates ACs into pre-trained multi-modal encoders (Sec.3.2). Subsequently, we detail the AC adaptation in Sec.3.3 and introduce DLFM for updating process in Sec.3.4. An overview of MDAA is provided in Figure 2.
3.1. Problem Definition and Notations
@@
To illustrate our method, we consider an audio-video classification task, which generalizes to other multi-modal scenarios without loss of generality. In MM-CTTA, the weight in pre-trained models is trained on a labeled source dataset in source domain Sr, where and represent the audio and video training data respectively. represents the corresponding one-hot label set. During the adaptation phase, for each timestamp in the target domain Tg, the model can only perform inference and update itself based on the unlabeled test dataset . The target domain varies with and the whole adaptation process can be represented as .
3.2. Source Model and Analytic Classifier Configuration
The multi-modal backbone follows a typical extraction–fusion architecture, where training data from different modalities are first processed by modality-specific feature encoders and subsequently fused through a fusion encoder. During this process, features from different encoders are projected into a unified higher-dimensional space to enhance representational capacity, and are then passed through a linear classifier to produce prediction labels, as illustrated in Figure 2.A. These linear classifiers are referred to as Analytic Classifiers (ACs). During the adaptation phase, both the feature encoders and the projection layers are kept frozen, while only the ACs are updated. Since features from different modalities are processed by ACs in a unified manner, we omit modality-specific distinctions for simplicity.
The weights for the ACs are optimized via ridge regression, initialized using a least squares objective, a strategy commonly employed in classification tasks (Zhuang et al., 2022; Fang et al., 2024; Zhuang et al., 2023; Yang and Hu, 2008). To address the issue of class imbalance, we reformulate the optimization objective as follows:
@@ 3.2. Source Model and Analytic Classifier Configuration
where is the number of classes in and is the number of samples in from category to which sample belongs. Following the ridge regression solution, the solution to optimization problem is given as
@@
where and . In addition to the classifier weights , a memory bank needs to be constructed during the training phase. Unlike other methods (Cao et al., 2023; Zhang et al., 2023; Xiong et al., 2024), our memory bank contains only two types of matrices, which can be represented as , where
@@
Both and are used to extract and preserve the learned knowledge from the source dataset, which cannot be accessed during adaptation. Therefore can be further rewritten as
3.3. Model Inference and Analytic Classifier Adaptation
During test time, MDAA determines the final prediction for each sample by selecting the output distribution with the highest maximum a posteriori (MAP) score across all ACs, as higher predictive probabilities generally reflect greater confidence (Boudiaf et al., 2022). Subsequently, DLFM is employed to further filter the samples used for model updates, which will be elaborated upon in the following section.
Since ACL performs global optimization, it considers not only the currently selected input sample at timestamp , but also all previously observed data, including and . The corresponding optimization problem for the weight is formulated as:
where denotes for soft pseudo-labels of filtered samples. We adopt soft labels for AC updates as they retain a richer representation of uncertainty compared to hard labels (Müller et al., 2019; Hinton et al., 2015). For each filtered sample, we sort and choose its top- class positions and assign weights . In this paper, we define the weights as , and construct the soft pseudo-label accordingly as:
Such reconsrtuction is proven to be useful, as shown in ablation study in subsection 4.3.